# FuxiTranyu: A Multilingual Large Language Model Trained with Balanced Data

### v2

@@ Abstract
Large language models (LLMs) have demonstrated prowess in a wide range of tasks. However, many LLMs exhibit significant performance discrepancies between high- and low-resource languages. To mitigate this challenge, we present FuxiTranyu, an open-source multilingual LLM, which is designed to satisfy the need of the research community for balanced and high-performing multilingual capabilities. FuxiTranyu-8B, the base model with 8 billion parameters, is trained from scratch on a meticulously balanced multilingual data repository that contains 600 billion tokens covering 43 natural languages and 16 programming languages. In addition to the base model, we also develop two instruction-tuned models: FuxiTranyu-8B-SFT that is fine-tuned on a diverse multilingual instruction dataset, and FuxiTranyu-8B-DPO that is further refined with DPO on a preference dataset for enhanced alignment ability. Extensive experiments on a wide range of multilingual benchmarks demonstrate the competitive performance of FuxiTranyu against existing multilingual LLMs, e.g., BLOOM-7B, PolyLM-13B, Llama-2-Chat-7B and Mistral-7B-Instruct. Interpretability analyses at both the neuron and representation level suggest that FuxiTranyu is able to learn consistent multilingual representations across different languages. To promote further research into multilingual LLMs and their working mechanisms, we release both the base and instruction-tuned FuxiTranyu models together with 58 pretraining checkpoints at HuggingFace [^1] https://huggingface.co/TJUNLP/FuxiTranyu-8B and Github. [^2] https://github.com/tjunlp-lab/FuxiTranyu
11footnotetext: Correspondence to: Deyi Xiong.
@@ 1 Introduction
A well-pretrained base model plays a pivotal role in facilitating research and applications of large language models. However, training a base LLM from scratch typically demands a substantial amount of data and significant computational resources, posing a barrier to the development of new LLMs. On the other hand, the majority of LLMs are usually tailored to specific languages such as English Touvron et al. (2023a, b) or Chinese Bai et al. (2023), neglecting the high demand for multilingual capabilities across multiple languages, especially low-resource languages. While certain LLMs, such as Mistral models Jiang et al. (2023a), demonstrate multilingual capabilities, their language coverage remains limited. This limitation significantly restricts the exploration of multilingualism in LLMs under the massive multilingual setting.
@@
Recent efforts have been dedicated towards mitigating such language-specific constraints through supervised fine-tuning, as exemplified by Okapi Lai et al. (2023). However, as highlighted by the alignment hypothesis in LIMA Zhou et al. (2024), the knowledge and capabilities of LLMs are predominantly derived from pre-training rather than supervised fine-tuning. Supervised fine-tuning primarily serves to align the behaviors of these models with instructions, which constitutes a sub-distribution of the pre-training data. Consequently, for LLMs whose pre-training data are dominated by a few languages, the effectiveness of supervised fine-tuning in enhancing their multilingual capabilities might be limited.
Other initiatives have focused on pre-training multilingual LLMs, such as BLOOM Scao et al. (2022a) and PolyLM Wei et al. (2023). Nevertheless, these efforts are hindered by their performance, which does not measure up to that of current trending LLMs. BLOOM suffers from outdated training data while PolyLM is undermined by imbalanced language distribution, with English data accounting for approximately 70% and Chinese for ~20%, potentially leading to insufficient learning of under-represented languages. Previous studies Xu et al. (2024) disclose three traits of multilingual LLMs caused by imbalanced language resources in pre-training: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, suggesting that multilingual LLMs could benefit from balanced data distribution across languages.
@@
In addition to the base model, we develop two instruction-tuned models, FuxiTranyu-8B-SFT that is fine-tuned on a collected high-quality multilingual instruction dataset, and FuxiTranyu-8B-DPO that is further tuned on preferences with DPO for enhanced alignment ability.
To evaluate multilingual capabilities of the FuxiTranyu models, we have conducted extensive evaluations across multiple domains, encompassing multilingual discriminative tasks such as multilingual ARC, HellaSwag, and MMLU Lai et al. (2023), XWinograd Muennighoff et al. (2022); Tikhonov and Ryabinin (2021), XCOPA Ponti et al. (2020), XStoryCloze Lin et al. (2021), and multilingual generative tasks including WMT and IWSLT translation benchmarks Bojar et al. (2016); Cettolo et al. (2017) and XL-Sum summarization benchmark Hasan et al. (2021). Our evaluations focus on knowledge, capability and alignment dimensions categorized by Guo et al. (2023). As detailed in Section 5, FuxiTranyu models have demonstrated superior performance on the multilingual ARC, HellaSwag, MMLU, XWinograd, XCOPA, and XStoryCloze compared to BLOOM-7B1 and PolyLM-13B. Furthermore, our two instruction-tuned models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, outperform Llama-2-Chat-7B, Mistral-7B-Instruct-v0.1, BLOOMZ-7B1, PolyLM-MultiAlpaca-13B on translation benchmarks. FuxiTranyu also achieves remarkable results on summarization.
To provide a deep understanding of the multilingual capabilities of FuxiTranyu models, we have conducted interpretability analyses from two distinct perspectives: neuron analysis and representation analysis, as detailed in Section 6. Analysis results indicate that FuxiTranyu-8B has learned more language-agnostic representations compared to BLOOM-7B1 Scao et al. (2022a), which can be attributed to the balanced distribution of our pre-training data. However, for languages with extremely limited resources and poor evaluation performance, such as Bengali and Tamil, FuxiTranyu-8B tends to allocate fewer neurons to process them. Additionally, different layers and components of FuxiTranyu-8B handle multilingual text differently, with deep layers being more language-specific and the importance of attention and MLP components varying across layers.
@@ 2 Related Work
The rapid advancement of LLMs has led to a surge in research on multilingual LLMs, aimed at supporting a broader range of languages and tasks. Training multilingual LLMs typically involves a multi-stage process, combining different approaches to enhance the model’s capabilities across multiple languages, either training a model from random initialization on massive multilingual data (e.g., BLOOM Scao et al. (2022a), OPT Zhang et al. (2022), PaLM Chowdhery et al. (2022), LLaMA Touvron et al. (2023a)) or building upon existing pretrained LLMs to reduce computational cost (e.g., X-Gen Vu et al. (2022), FinGPT Luukkonen et al. (2023), Cabrita Larcher et al. (2023), Sabia Almeida et al. (2024)). While these methods have made significant strides in bridging the gap between high- and low-resource languages, challenges still remain.
@@
From-scratch pre-training often struggles with the curse of multilinguality, where adding more languages can lead to performance degradation for low-resource languages. Continual pre-training, while more efficient, suffers from catastrophic forgetting, where models forget previously learned knowledge. Supervised fine-tuning (SFT) often leverages multilingual instruction data or incorporates translation tasks to address data scarcity Shen et al. (2023a); Lai et al. (2023); Wang et al. (2022). However, both methods rely heavily on high-quality, diverse datasets, which are often limited for many languages. Reinforcement Learning from Human Feedback (RLHF) is increasingly used to align models with human preferences Shen et al. (2023b). In multilingual LLMs, multilingual RLHF data are used to train multilingual reward models Chen et al. (2024). However, RLHF typically relies on human-annotated data, which can be expensive and time-consuming to collect, especially for under-resourced languages. Downstream fine-tuning involves either tuning all parameters on downstream tasks Rosenbaum et al. (2022); Yang et al. (2023) or employing parameter-efficient fine-tuning methods to reduce costs Tu et al. (2024); Whitehouse et al. (2023). While these methods can achieve impressive performance, they can also be computationally expensive and may not generalize well to unseen tasks or languages.
Recent years have witnessed that prominent MLLMs have been developed, each with specific training methodologies and strengths. These include BLOOM (176B parameters, open-source, over 46 languages), LLaMA (65B parameters, efficient architecture), PaLM (540B parameters, wide benchmark success), OPT (175B parameters, open-source), Qwen (14B parameters, strong benchmark performance ), Mistral (7B parameters, open-source, competitive performance ), and Orion-14B (14B parameters, diverse data of 2.5T tokens, data scheduling strategy). While these models have achieved impressive results, future work should focus on addressing the limitations of existing approaches. We strongly suggest that efforts should be made to develop more robust and efficient training methods and strategies that address the curse of multilinguality, mitigate catastrophic forgetting, alleviate data imbalance, and minimize reliance on expensive annotated data, especially for low-resource languages.
ISO-931 Language Language Family ISO-931 Language Language Family
ar Arabic Afro-Asiatic ky Kyrgyz Turkic
bg Bulgarian Indo-European lo Lao Kra-Dai
bn Bengali Indo-European ms Malay Austronesian
ca Catalan Indo-European my Burmese Sino-Tibetan
cs Czech Indo-European nl Dutch Indo-European
de German Indo-European pl Polish Indo-European
el Greek Indo-European pt Portuguese Indo-European
en English Indo-European ro Romanian Indo-European
es Spanish Indo-European ru Russian Indo-European
fa Persian Indo-European sv Swedish Indo-European
fi Finnish Uralic ta Tamil Dravidian
fr French Indo-European tg Tajik Indo-European
he Hebrew Afro-Asiatic th Thai Kra-Dai
hi Hindi Indo-European tk Turkmen Turkic
hu Hungarian Indo-European tl Filipino Austronesian
id Indonesia Austronesian tr Turkish Turkic
it Italian Indo-European uk Ukrainian Indo-European
ja Japanese Japanic ur Urdu Indo-European
kk Kazakh Turkic uz Uzbek Turkic
km Khmer Austroasiatic vi Vietnamese Austroasiatic
ko Korean Koreanic zh Chinese Sino-Tibetan
ku Kurdish Indo-European
Table 2: The list of 43 natural languages supported by FuxiTranyu.
Language Size (GB) Ratio (%) Language Size (GB) Ratio (%)
Java 96 17.94 Go 26 4.86
JavaScript 70 13.08 SQL 11 2.06
Python 63 11.77 Rust 9.1 1.70
PHP 59 11.02 Ruby 7.9 1.48
C 53 9.90 Scala 5.1 0.95
C++ 52 9.72 Lua 3.0 0.56
C# 48 8,97 Assembly 1.6 0.30
TypeScript 29 5.42 Visual Basic 1.5 0.28
Table 3: The list of 16 programming languages covered in FuxiTranyu, including the sizes and ratios of each language.
@@ 3 Pretraining
We present the strategy we used to determine which languages should be supported by FuxiTranyu series of models in Section 3.1. After that, we elaborate the sources and domains of our pre-training data, and the efforts we have made in the pre-processing stage in Section 3.2. Next, we discuss the details of our tokenizer training in Section 3.3 and the details of our FuxiTranyu architecture in Section 3.4. Finally, we present the pre-training settings in Section 3.5.
3.1 Supported Languages in FuxiTranyu
@@ 3.2 Data Collection
A significant portion of our multilingual data comprises web documents, as they provide a vast amount of data for pre-training, akin to other open-sourced LLMs Touvron et al. (2023a); Bai et al. (2023); Cai et al. (2024); Young et al. (2024). We opt to utilize CulturaX Nguyen et al. (2023), a filtered subset of OSCAR Ortiz Su’arez et al. (2020); Suárez et al. (2019) (itself a subset of Common Crawl) and mC4 Raffel et al. (2020) datasets. To enhance the quality and diversity of our pre-training corpora, we further collect data from various sources such as ROOTS Laurençon et al. (2022), MultiUN Eisele and Chen (2010); Chen and Eisele (2012), and OpenSubtitles Lison and Tiedemann (2016). We primarily select documents in languages included in our language list. We further include data sourced from encyclopedias and reports. Inspired by the Phi series models Gunasekar et al. (2023), which leverage high-quality data from textbooks to achieve remarkable performance, we also integrate books and articles data into our final data mixture. Approximately 500GB of articles data have been gathered from Semantic Scholar (S2ORC) Lo et al. (2020), and around 10GB of Chinese books data sourced from Fudan Cbook dataset. [^5] https://github.com/FudanNLPLAB/CBook-150K
@@
Multilingual book data are obtained from Project Gutenberg based on the provided language identity, although it constitutes a small portion of our final corpora. Additionally, we collect 535GB of code data from open-source datasets. The primary source is Starcoderdata, [^6] https://huggingface.co/datasets/bigcode/starcoderdata a subset of the Stack dataset Kocetkov et al. (2022) used to train the StarCoder model Li et al. (2023). We also include a subset of Github code from the RedPajama dataset. [^7] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T
@@
At the filtering stage, we primarily employ three different filtering methods, aligning with previous works Scao et al. (2022a); Almazrouei et al. (2023); Bai et al. (2023); Young et al. (2024). The initial filtering phase incorporates heuristic rules to exclude undesired documents. This involves filtering out documents containing URLs or words listed in blacklists, such as stop words or flagged words. Subsequently, we filter documents based on statistical information, including the ratio or number of repeated n-gram characters or words, as well as the document length. Following this, we apply a learned quality filter method based on specific metrics, such as perplexity. In line with the approach taken in BLOOM Scao et al. (2022a), we utilize KenLM Heafield (2011) to compute the perplexity of the documents and subsequently filter out those surpassing the pre-defined threshold.
@@
Upon completion of the quality filter stage, significant efforts are dedicated to data deduplication, as previous studies have emphasized its importance for LLM performance Lee et al. (2022). We employ fuzzy-match deduplication using the MinHash algorithm. However, due to the memory-intensive nature of deduplication, processing the entire dataset at once on a server with limited memory is unfeasible. Yet, processing only a portion of the data will not achieve complete deduplication. To address this challenge, we apply a strategy of multi-turn micro-deduplication. We first split large documents into multiple chunks and maintain a chunk pool. In each turn, we randomly select chunks from the pool and perform deduplication among these chunks. Once processed, these collected chunks are randomly split into multiple chunks and reintegrated into the chunk pool. This procedure is repeated multiple times until the number of filtered-out documents is less than 1%. In practice, we employ multi-turn deduplication primarily for high-resource languages. For low-resource languages, the entire dataset could fit into memory at once due to the limited amount of pre-training data. In the case of code data, we also utilize the MinHash algorithm for data deduplication. Specifically, we leverage the implementation from the bigcode project. [^8] https://github.com/bigcode-project/bigcode-dataset/blob/main/near_deduplication/minhash_deduplication.py
3.3 Tokenization
Following the approach used in BLOOM, we also evaluate the performance of our tokenizer using the fertility metric. To assess its efficacy, we conduct a comparative analysis with the Llama-2 and BLOOM tokenizers. This evaluation involves computing fertility on the same set of documents across different languages. Results are presented in Figure 2, which indicate that the FuxiTranyu tokenizer is more efficient than the others in most languages. Based on our evaluations and interpretability analysis, we believe that the fertility of the tokenizer positively correlates with the model’s performance on specific languages. In the fertility test, we observe that Bengali (bn), Hindi (hi), and Tamil (ta) exhibit high fertility, indicating lower tokenization efficiency in these languages compared to others. Consequently, the performance and importance of neurons of these languages in our base model are also suboptimal. Further details are discussed in Section 6.1.2.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated prowess in a wide range of tasks. However, many LLMs exhibit significant performance discrepancies between high- and low-resource languages. To mitigate this challenge, we present FuxiTranyu, an open-source multilingual LLM, which is designed to satisfy the need of the research community for balanced and high-performing multilingual capabilities. The base model, FuxiTranyu-8B, features 8 billion parameters and is trained from scratch on meticulously balanced multilingual data that contains 600 billion tokens covering 43 natural languages and 16 programming languages. We also develop two instruction-tuned models: FuxiTranyu-8B-SFT which is fine-tuned on a diverse multilingual instruction dataset, and FuxiTranyu-8B-DPO which is further refined with DPO on a preference dataset for enhanced alignment ability. Extensive experiments on a wide range of multilingual benchmarks demonstrate the competitive performance of FuxiTranyu against existing multilingual LLMs, e.g., BLOOM-7B, PolyLM-13B, and Mistral-7B-Instruct. Both neuron and representation interpretability analyses reveal that FuxiTranyu achieves consistent multilingual representations across languages. To promote further research into multilingual LLMs, we release both the base and instruction-tuned FuxiTranyu models together with 58 pre-training checkpoints at HuggingFace [^1] https://huggingface.co/TJUNLP/FuxiTranyu-8B and Github. [^2] https://github.com/tjunlp-lab/FuxiTranyu
11footnotetext: Corresponding author.
@@ 1 Introduction
A well-pretrained base model is crucial for facilitating research and applications of large language models. However, training a base LLM from scratch typically demands a substantial amount of data and significant computational resources, posing a barrier to the development of new LLMs. The majority of LLMs are usually tailored to specific languages such as English Touvron et al. (2023a, b) or Chinese Bai et al. (2023), neglecting the growing demand for multilingual capabilities, especially from low-resource languages. While certain LLMs like Mistral models Jiang et al. (2023a) demonstrate multilingual capabilities, their coverage is limited, restricting the exploration in massively multilingual settings.
@@
Recent efforts have been dedicated towards mitigating such language-specific constraints through supervised fine-tuning, as exemplified by Okapi Lai et al. (2023). However, as highlighted by the alignment hypothesis in LIMA Zhou et al. (2024), the knowledge of LLMs is predominantly derived from pre-training, while supervised fine-tuning primarily aligns model behavior to instructions, which is a narrow subset of the pre-training data. This makes fine-tuning less effective for boosting multilingual abilities when pre-training is dominated by a few languages.
Other initiatives have focused on pre-training multilingual LLMs, such as BLOOM Scao et al. (2022a) and PolyLM Wei et al. (2023). Nevertheless, these efforts are hindered by their performance, which does not measure up to that of current trending LLMs. BLOOM suffers from outdated training data, while PolyLM is undermined by imbalanced language distribution, with English data accounting for approximately 70% and Chinese for ~20%, potentially leading to insufficient learning of under-represented languages. Previous studies Xu et al. (2024) disclose three traits of multilingual LLMs caused by imbalanced language resources: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional transfer between high- and low-resource languages, emphasizing the need for balanced data distribution.
@@
In addition to the base model, we develop two instruction-tuned models: FuxiTranyu-8B-SFT, fine-tuned on a collected high-quality multilingual instruction dataset, and FuxiTranyu-8B-DPO, further tuned on preferences with DPO for enhanced alignment ability.
Our evaluations focus on knowledge, capability and alignment dimensions categorized by Guo et al. (2023). Evaluation results on multilingual discriminative tasks such as multilingual ARC, HellaSwag, and MMLU Lai et al. (2023), XWinograd Muennighoff et al. (2022); Tikhonov and Ryabinin (2021), XCOPA Ponti et al. (2020), XStoryCloze Lin et al. (2021), and multilingual generative tasks including WMT and IWSLT translation benchmarks Bojar et al. (2016); Cettolo et al. (2017) and XL-Sum summarization benchmark Hasan et al. (2021), demonstrate FuxiTranyu’s superior performance compared to BLOOM-7B1 and PolyLM-13B, as detailed in Section 5. The instruction-tuned models, FuxiTranyu-8B-SFT and FuxiTranyu-8B-DPO, also outperform Llama-2-Chat-7B, Mistral-7B-Instruct-v0.1, BLOOMZ-7B1, PolyLM-MultiAlpaca-13B on translation and summarization benchmarks.
To further understand the multilingual capabilities of FuxiTranyu models, we have conducted neuron- and representation-level analysis, revealing that FuxiTranyu-8B learns more language-agnostic representations compared to BLOOM-7B1 Scao et al. (2022a), which can be attributed to the balanced pre-training data. However, languages with extremely limited resources, such as Bengali and Tamil, are allocated with fewer neurons. Additionally, different layers and components of FuxiTranyu-8B handle multilingual text differently, with deep layers being more language-specific and the importance of attention and MLP components varying across layers.
@@ 2 Related Work
Recent advanced LLMs Touvron et al. (2023a, b); Dubey et al. (2024); Bai et al. (2023); Yang et al. (2024); Young et al. (2024); Jiang et al. (2023a); Team et al. (2024a, b) have excelled in NLP and cross-modal tasks, sparking increased research on multilingual LLMs Scao et al. (2022a); Chowdhery et al. (2022); Wei et al. (2023), which aim at broader language support. There are three main approaches to building multilingual LLMs: pre-training from scratch, continual pre-training, and post-training (e.g., supervised fine-tuning and reinforcement learning from human feedback).
Pre-training from scratch, like PaLM 1&2 Chowdhery et al. (2022); Anil et al. (2023), BLOOM Scao et al. (2022a), and PolyLM Wei et al. (2023), leverages extensive training corpora from diverse sources, enabling the incorporation of new knowledge. However, pre-training poses a variety of challenges, such as the need for vast computing resources, which can hinder the development of new multilingual LLMs. Additionally, it also suffers from the curse of multilinguality Conneau et al. (2019); Chai et al. (2022); Dubey et al. (2024); Gurgurov et al. (2024), where the performance of individual languages deteriorates as the number of languages increases. On the other hand, continual pre-training, like Aurora-M Nakamura et al. (2024), LLaMAX Lu et al. (2024), is more efficient but risks catastrophic forgetting of previously learned knowledge.
@@
Supervised fine-tuning (SFT) often leverages multilingual instruction data or incorporates translation tasks to address data scarcity Shen et al. (2023a); Lai et al. (2023); Wang et al. (2022). However, both continual pre-training and SFT rely heavily on high-quality, diverse datasets, which are often limited to many languages. Reinforcement Learning from Human Feedback (RLHF) is increasingly used to align models with human preferences Shen et al. (2023b). In multilingual LLMs, multilingual RLHF data are used to train multilingual reward models Chen et al. (2024). However, RLHF typically relies on human-annotated data, which can be expensive and time-consuming to collect, especially for under-resourced languages. While these methods can achieve impressive performance, they can also be computationally expensive and may not generalize well to unseen languages.
@@ 3 Pretraining
We elaborate on the sources and domains of our pre-training data and the efforts we have made in the pre-processing stage in Section 3.1. Next, we discuss the details of our FuxiTranyu architecture in Section 3.2. We present the strategy we used to determine which languages should be supported by the FuxiTranyu series of models in Appendix A, the details of our tokenizer training in Appendix B, and the pre-training settings in Section C.1.
@@ 3.1 Data Collection
A significant portion of our multilingual data comprises web documents, a common approach in open-sourced LLMs Touvron et al. (2023a); Bai et al. (2023); Cai et al. (2024); Young et al. (2024). We opt to utilize CulturaX Nguyen et al. (2023), a filtered subset of OSCAR Ortiz Su’arez et al. (2020); Suárez et al. (2019) (itself a subset of Common Crawl) and mC4 Raffel et al. (2020) datasets. To improve quality and diversity, we supplement these with data from ROOTS Laurençon et al. (2022), MultiUN Eisele and Chen (2010); Chen and Eisele (2012), and OpenSubtitles Lison and Tiedemann (2016), focusing on languages in our language list. Additionally, we incorporate data from encyclopedias, reports, books, and articles, drawing inspiration from Phi series models Gunasekar et al. (2023) that achieve strong results using high-quality textbooks. We have collected approximately 500GB of article data from Semantic Scholar (S2ORC) Lo et al. (2020), and around 10GB of Chinese books from the Fudan Cbook dataset. [^4] https://github.com/FudanNLPLAB/CBook-150K We also source multilingual book data from Project Gutenberg, though it forms a small portion of the final corpus.
@@
Additionally, we collect 535GB of code data from open-source datasets, primarily from Starcoder data, [^5] https://huggingface.co/datasets/bigcode/starcoderdata a subset of the Stack dataset Kocetkov et al. (2022) used to train the StarCoder model Li et al. (2023). We also include a subset of Github code from the RedPajama dataset. [^6] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T
@@
At the filtering stage, we primarily employ three methods similar to prior works Scao et al. (2022a); Almazrouei et al. (2023); Bai et al. (2023); Young et al. (2024). The initial filtering phase uses heuristic rules to exclude undesired documents. This involves filtering out documents containing blacklisted URLs or words, such as stop words or flagged words. Subsequently, we filter documents based on statistical information, including the ratio or the number of repeated n-gram characters or words, as well as the document length. Following this, we apply a learned quality filter method based on specific metrics, such as perplexity. In line with the approach taken in BLOOM Scao et al. (2022a), we utilize KenLM Heafield (2011) to compute the perplexity of the documents and subsequently filter out those exceeding a predefined threshold.
@@
Upon completion of the quality filter stage, significant efforts are dedicated to data deduplication, as previous studies have emphasized its importance for LLM performance Lee et al. (2022). We employ MinHash for fuzzy-match deduplication. However, due to the memory-intensive nature of deduplication, processing the entire dataset at once on a server with limited memory is unfeasible. Yet, processing only a portion of the data will not achieve complete deduplication. To address this challenge, we apply a strategy of multi-turn micro-deduplication. We split large documents into chunks and store them in a chunk pool. In each turn, we randomly select chunks from the pool, assemble them back into documents, and perform deduplication on these assembled documents. After processing, the deduplicated documents are again split into chunks and reintegrated into the pool. This process is repeated multiple times until the number of filtered-out documents drops below 1%. This approach is used for high-resource languages, while low-resource languages are processed in memory due to their smaller dataset size. In the case of code data, we also utilize the MinHash algorithm for data deduplication. Specifically, we leverage the implementation from the bigcode project. [^7] https://github.com/bigcode-project/bigcode-dataset/blob/main/near_deduplication/minhash_deduplication.py